# Lambda-Skip Connections: the architectural component that prevents rank collapse

### v1

@@ Abstract
Rank collapse, a phenomenon where embedding vectors in sequence models rapidly converge to a uniform token or equilibrium state, has recently gained attention in the deep learning literature. This phenomenon leads to reduced expressivity and potential training instabilities due to vanishing gradients. Empirical evidence suggests that architectural components like skip connections, LayerNorm, and MultiLayer Perceptrons (MLPs) play critical roles in mitigating rank collapse. While this issue is well-documented for transformers, alternative sequence models, such as State Space Models (SSMs), which have recently gained prominence, have not been thoroughly examined for similar vulnerabilities. This paper extends the theory of rank collapse from transformers to SSMs using a unifying framework that captures both architectures. We study how a parametrized version of the classic skip connection component, which we call lambda-skip connections, provides guarantees for rank collapse prevention. Through analytical results, we present a sufficient condition to guarantee prevention of rank collapse across all the aforementioned architectures. We also study the necessity of this condition via ablation studies and analytical examples. To our knowledge, this is the first study that provides a general guarantee to prevent rank collapse, and that investigates rank collapse in the context of SSMs, offering valuable understanding for both theoreticians and practitioners. Finally, we validate our findings with experiments demonstrating the crucial role of architectural components such as skip connections and gating mechanisms in preventing rank collapse.{NoHyper} Preprint. Under Review.
@@ 1 Introduction
The phenomenon of rank collapse has been recently reported in the deep learning literature Dong et al. [2023], Noci et al. [2022], Shi et al. [2022], He et al. [2023], Geshkovski et al. [2024], Wu et al. [2024a, b]. Rank collapse is a phenomenon by which embedding vectors exhibit a fast convergence rate into a uniform token, or equilibrium embedding vector. Various explanations why this phenomenon arises have been proposed: Dong et al. [2023] proposed a path decomposition argument and suggests that one of the reasons on why rank collapse happens is that the attention matrix is row stochastic. Geshkovski et al. [2024] instead proposes an analysis of rank collapse from a statistical physics perspective by treating transformers as mean-field interacting particle systems and observes that the resulting system converges to a metastable state corresponding to a Dirac delta measure. In practice, this convergence to a uniform token reduces model expressivity Daneshmand et al. [2020], Dong et al. [2023], Noci et al. [2022], Wu et al. [2024b] since a low rank in the output matrix means that the model is unable to capture complex relationships between tokens. Furthermore, this phenomenon can also cause vanishing gradients during training [Noci et al., 2022].
@@
Sequence models alternative to transformers have recently gained widespread attention. In particular, State Space Models (SSMs) such as S4 Gu et al. [2022b], S5 Smith et al. [2023], H3 Fu et al. [2023] and Mamba Dao and Gu [2024], Gu and Dao [2024] have been very prominent, where attention was replaced by SSM blocks, which operate in combination with other components such as Gating Mechanisms (Hochreiter and Schmidhuber [1997], Chung et al. [2014]), MLPs (Popescu et al. [2009] and skip connections (He et al. [2015]). Furthermore, comparisons between these architectures have been discussed in [Sieber et al., 2024, Hè and Kabic, 2023, Tiezzi et al., 2024], and a larger framework where both the transformer and SSMs are seen as different realizations of the same general model has recently been proposed (Dao and Gu [2024]). However, whether they suffer from similar weaknesses, and whether rank collapse happens more broadly in these architectures, is still an open question.
In this work, we focus on the specific issue of rank collapse and explore how slight modifications in the architecture, referred to as lambda-skip connections, can help to prevent it. Although parametrized versions of skip connections have been explored in the past (see e.g. He et al. [2016],Srivastava et al. [2015]), their impact in providing guarantees to prevent rank collapse has not been studied. Moreover, here we develop a general theory of rank collapse using the unifying framework for sequence models [Dao and Gu, 2024], comprising both transformers and SSM architectures. To the best of our knowledge, the rank collapse phenomenon in the SSM setting has not yet been studied. Since rank collapse can be a very serious issue for practitioners, investigating broadly when rank collapse occurs and designing robust architectures that prevent rank collapse is of central interest not only to theorists but also to experimentalists. Our contributions can be outlined as follows:
  • We extend the theory on rank collapse originally developed for transfomers to SSMs by using the recently proposed framework unifying these two architecture types [Dao and Gu, 2024].
@@
  • We also empirically show the role of gating mechanisms in preventing rank collapse. To the best of our knowledge, we are the first to make this connection between gating mechanisms, which were originally constructed for improving memory capabilities of the models (Hochreiter and Schmidhuber [1997]), and rank collapse.
@@ 2 Related Works
Rank collapse is the phenomenon in which the rank of the output layer of the network collapses to 1 (Dong et al. [2023]). To mathematically capture this, measuring the rank is not ideal since for every finite amount of layers, the output matrix will always be full rank with probability 1. This led to the development of an alternative metric to measure rank collapse Dong et al. [2023], which intuitively evaluates how “ill-conditioned” is the matrix of the output layer, i.e., how close it is to being rank $1$. In Dong et al. [2023], this metric is defined by computing how far the columns of the matrix are from the mean of all columns. It was shown in e.g. Dong et al. [2023], Wu et al. [2024a] that this metric quickly decays to 0 as the depth of the model increases. Hence, rank collapse is an intrinsic issue of deep-learning models. Rank collapse has mainly been studied for transformers (Vaswani et al. [2023]), Graph Neural Networks (Zhou et al. [2021]) and ReLU networks (Agarap [2019]), but not SSMs. The work in Daneshmand et al. [2020] studied rank collapse in randomly initialized linear and ReLU networks and proved that Batch Normalization can be helpful to mitigate the phenomenon. In Dong et al. [2023] the authors show that in self-attention-only transformers, rank collapse occurs doubly exponentially, i.e. with rate $O(a^{b^{n}})$ for some $a,b\in\mathbb{R}$. In the same paper, the authors propose for the first time that skip connections and MLPs (Popescu et al. [2009]) are helpful in preventing rank collapse from happening. Moreover, the work in Noci et al. [2022] shows that rank collapse is not only an issue during inference but also hinders training due to vanishing gradient problems at initialization. In Wu et al. [2024b] the authors show that rank collapse also happens in graph attention networks (in this setting the phenomenon is also known as oversmoothing), also causing the model to lose expressive power. Finally, the work in Wu et al. [2024a] studies the effect of LayerNorm on rank collapse in transformers, by proving that both of these components can help preventing or mitigating rank collapse.
@@ 3.1 Sequence Model Deep-Learning Architecture
The main mechanism differs between attention in transformers, and a recurrent block in SSMs. However, as shown in Dao and Gu [2024], both models can be seen as part of a general framework
@@
where $A^{(k)}_{t}\in\mathbb{R}^{dH\times dH},B^{(k)}_{t}\in\mathbb{R}^{dH\times 1},C^{(k)}_{t}\in\mathbb{R}^{1\times dH},D^{(k)}_{t}\in\mathbb{R}^{1\times 1}$. The first versions of the SSM blocks consisted of Linear Time Invariant (LTI) representations, i.e., $A_{t}=A$, $B_{t}=B$, $C_{t}=C$ and $D_{t}=D$ for all time steps $t$ [Gu et al., 2022b, a, Smith et al., 2023]. Recently, in order to improve the expressivity and the performance of such models, selective State Space Models have been proposed [Gu and Dao, 2024, Dao and Gu, 2024]. In these models, matrices $A_{t}$, $B_{t}$, and $C_{t}$ are input-dependent. We refer to Appendix A.1 for a detailed explanation of attention and the recurrent block.
@@ 4.1 Lambda-Skip Connection: sufficient to prevent rank collapse
$b:=\frac{1}{a^{K}}\frac{2\lambda NS^{2}C_{M}}{\lambda^{2}c^{2}-aS^{2}(C_{M}+\lambda)^{2}},$
[…]
@@
The key part of the proof is to show the following relationship between the rank collapse measure at two consecutive layers: $\mu(Y^{(k+1)})^{2}\geq\frac{1}{S^{2}(C_{M}+\lambda)^{2}}\left(\lambda^{2}c^{2}\mu(Y^{(k)})^{2}-2\lambda NS^{2}C_{M}\right)$, which is done by algebraic manipulation and upper bounds on $||Y^{(k+1)}||_{F}$. The claim then follows by imposing the right-hand side above greater of equal to $a\mu(Y^{(k)})^{2}$, rearranging the terms and imposing the condition on $\lambda$ to guarantee the feasibility of this. The full proof can be found in Appendix A.4. ∎
We note that this result is very general, since it holds for any model that can be expressed in the form of equation 6, independent of its specific implementation or parametrization. Moreover, the presence of LayerNorm implies that the bound provided in equation 7 does not depend on the input even when the matrix $M$ is input-dependent. This is because the input is normalized before being fed through $M$. Specifically, the constant $C_{M}$ is only dependent on the weights and on the sequence length  [^3] In transformers $C_{M}=\sqrt{N}$, in structured LTI SSMs $C_{M}\geq\sup_{k}||A^{(k)}||\ ||B^{(k)}||\ ||C^{(k)}||_{F}$ and in selective SSMs $C_{M}\geq\sup_{k}||W_{B}^{(k)}||\ ||W_{C}^{(k)}||_{F}$ (since in Selective SSMs it is common practice to choose $A_{t}^{(k)}$ with eigenvalues smaller than one).
Theorem 4.1 also provides an important intuition for the choice of $\lambda$. In particular, we note that to guarantee prevention of rank collapse, the ideal choice is $a=1$. However, this choice will be mediated by the values of $c$ and $S$ in the different architectures. In particular, it is easy to see that the only way to guarantee a solution to 7 is by having $c^{2}-aS^{2}>0$, for which $\lambda$ needs to satisfy $|\lambda|>\frac{aC_{M}S^{2}+\sqrt{ac^{2}C_{M}^{2}S^{2}}}{c^{2}-aS^{2}}$.
@@ 4.2.1 Rank collapse occurs without skip connections
We refer to Appendix A.5 for a rigorous presentation of Theorem 4.2. The main takeaway from this result is that, under appropriate conditions on the architecture and the input sequence, removing skip connections and only using attention layers and LayerNorm can cause an exponentially fast convergence to a rank-$1$ matrix. This highlights the importance of skip connections to mitigate rank collapse. It is also interesting to compare the result in Theorem 4.2 to the setting of self-attention only transformers, i.e. removing both LayerNorm and skip connections, reported in Dong et al. [2023]. In particular, under certain assumptions on the attention matrix and the input sequence, rank collapse occurs at a doubly exponential rate. We report the full result for completeness in Theorem A.3 and Corollary A.3.1 (where we adapted the Theorem to the rank collapse measure provided in the problem statement).
@@ 4.2.2 Rank Collapse can still occur with skip connections under specific choices of $\lambda$: examples
Here we show analytical examples of how, for different architectures, rank collapse occurs in the presence of lambda-skip connections, if $\lambda$ does not satisfy the condition in Theorem 4.1.
Structured LTI SSMs.
Proposition 4.3.1.
Proof.
The result follows by calculating $Y^{(k)}$ inductively and by showing that the corresponding $\mu(Y^{(k)})$ decays to 0. See Appendix A.10 for the full proof. ∎
@@
Consider the system $A_{t}=A=\alpha I$ (for simplicity, we consider $\alpha$ not to be input dependent here), $B=YW_{B}$ and $C=YW_{C}$ input dependent (where $y_{t}$ is the $t$-th row of $Y$) with LayerNorm applied to the output of each layer. Once again, we let $N=2$ and $d=2$. We have $M=W_{A}\odot\left(YW_{C}W_{B}^{\top}Y^{\top}\right)$ where $W_{A}=\begin{pmatrix}1&0\\ \alpha&1\\ \end{pmatrix}$. By choosing $\alpha=1$ and $W_{B}=W_{C}=I$, we get that $M=\begin{pmatrix}1&0\\ y_{1}^{\top}y_{2}&1\end{pmatrix}$ since the input to the layer is normalized by LayerNorm. Let us denote this system as [Sys-2].

### v3

@@ Abstract
Rank collapse, a phenomenon where embedding vectors in sequence models rapidly converge to a uniform token or equilibrium state, has recently gained attention in the deep learning literature. This phenomenon leads to reduced expressivity and potential training instabilities due to vanishing gradients. Empirical evidence suggests that architectural components like skip connections, LayerNorm, and MultiLayer Perceptrons (MLPs) play critical roles in mitigating rank collapse. While this issue is well-documented for transformers, alternative sequence models, such as State Space Models (SSMs), which have recently gained prominence, have not been thoroughly examined for similar vulnerabilities. This paper extends the theory of rank collapse from transformers to SSMs using a unifying framework that captures both architectures. We study how a parametrized version of the classic skip connection component, which we call lambda-skip connections, provides guarantees for rank collapse prevention. Through analytical results, we present a sufficient condition to guarantee prevention of rank collapse across all the aforementioned architectures. We also study the necessity of this condition via ablation studies and analytical examples. To our knowledge, this is the first study that provides a general guarantee to prevent rank collapse, and that investigates rank collapse in the context of SSMs, offering valuable understanding for both theoreticians and practitioners. Finally, we validate our findings with experiments demonstrating the crucial role of architectural components such as skip connections and gating mechanisms in preventing rank collapse.
@@ 1 Introduction
The phenomenon of rank collapse has been recently reported in the deep learning literature Dong et al. [2023], Noci et al. [2022], Shi et al. [2022], He et al. [2023], Geshkovski et al. [2024], Wu et al. [2024a, b]. Rank collapse is a phenomenon by which embedding vectors exhibit a fast convergence rate into a uniform token, or equilibrium embedding vector. Various explanations why this phenomenon arises have been proposed: Dong et al. [2023] proposed a path decomposition argument and suggests that one of the reasons on why rank collapse happens is that the attention matrix is row stochastic. Geshkovski et al. [2024] instead proposes an analysis of rank collapse from a statistical physics perspective by treating transformers as mean-field interacting particle systems and observes that the resulting system converges to a metastable state corresponding to a Dirac delta measure. In practice, this convergence to a uniform token reduces model expressivity Daneshmand et al. [2020], Dong et al. [2023], Noci et al. [2022], Wu et al. [2024b] since a low rank in the output matrix means that the model is unable to capture complex relationships between tokens, also potentially causing vanishing gradients during training [Noci et al., 2022]. Therefore, gaining a deeper understanding of this phenomenon—its origins, conditions for occurrence, and potential prevention strategies—is essential for designing models that are more robust, stable, and expressive.
@@
Sequence models alternative to transformers have recently gained widespread attention. In particular, State Space Models (SSMs) such as S4 Gu et al. [2022b], S5 Smith et al. [2023], H3 Fu et al. [2023] and Mamba Dao and Gu [2024], Gu and Dao [2024] have been very prominent, where attention was replaced by SSM blocks, which operate in combination with other components such as Gating Mechanisms (Hochreiter and Schmidhuber [1997], Chung et al. [2014]), MLPs (Popescu et al. [2009] and skip connections (He et al. [2015]). Furthermore, comparisons between these architectures have been discussed in [Sieber et al., 2024, Hè and Kabic, 2023, Tiezzi et al., 2024], and a larger framework where both the transformer and SSMs are seen as different realizations of the same general model has recently been proposed (Ali et al. [2024], Dao and Gu [2024]). However, whether they suffer from similar weaknesses, and whether rank collapse happens more broadly in these architectures, is still an open question.
In this work, we focus on the specific issue of rank collapse and explore how slight modifications in the architecture, referred to as lambda-skip connections, can help to prevent it. Although parametrized versions of skip connections have been explored in the past (see e.g. He et al. [2016],Srivastava et al. [2015]), their impact in providing guarantees to prevent rank collapse has not been studied. Moreover, here we develop a general theory of rank collapse using the unifying framework for sequence models presented in Ali et al. [2024] and [Dao and Gu, 2024], comprising both transformers and SSM architectures. To the best of our knowledge, the rank collapse phenomenon in the SSM setting has not yet been studied and we are the first to provide a general lower bound for rank collapse which holds for both Transformers and SSMs. Additionally, we offer a mechanistic solution to prevent rank collapse by introducing a skip strength parameter, which comes with minimal computational overhead while enhancing the model’s expressivity and stability. Since rank collapse can be a very serious issue for practitioners, investigating broadly when rank collapse occurs and designing robust architectures that prevent rank collapse is of central interest not only to theorists but also to experimentalists. Our contributions can be outlined as follows:
  • We extend the theory on rank collapse originally developed for transfomers to SSMs by using the recently proposed framework unifying these two architecture types [Ali et al., 2024, Dao and Gu, 2024].
@@
  • We also empirically show the role of gating mechanisms in preventing rank collapse. To the best of our knowledge, we are the first to make this connection between gating mechanisms, which were originally constructed for improving memory capabilities of the models [Hochreiter and Schmidhuber, 1997], and rank collapse.
@@ 2 Related Works
Rank collapse is the phenomenon in which the rank of the output layer of the network collapses to 1 (Dong et al. [2023]). To mathematically capture this, measuring the rank is not ideal since for every finite amount of layers, the output matrix will always be full rank with probability 1. This led to the development of an alternative metric to measure rank collapse Dong et al. [2023], which intuitively evaluates how “ill-conditioned” is the matrix of the output layer, i.e., how close it is to being rank $1$. In Dong et al. [2023], this metric is defined by computing how far the columns of the matrix are from the mean of all columns. It was shown in e.g. Dong et al. [2023], Wu et al. [2024a] that this metric quickly decays to 0 as the depth of the model increases. Hence, rank collapse is an intrinsic issue of deep-learning models. Rank collapse has mainly been studied for transformers (Vaswani et al. [2023]), Graph Neural Networks (Zhou et al. [2021]) and ReLU networks (Agarap [2019]), but not SSMs. The work in Daneshmand et al. [2020] studied rank collapse in randomly initialized linear and ReLU networks and proved that Batch Normalization can be helpful to mitigate the phenomenon. In Dong et al. [2023] the authors show that in self-attention-only transformers, rank collapse occurs doubly exponentially, i.e. with rate $O(a^{b^{n}})$ for some $a,b\in\mathbb{R}$. In the same paper, the authors propose for the first time that skip connections and MLPs (Popescu et al. [2009]) are helpful in preventing rank collapse from happening. Moreover, the work in Noci et al. [2022] shows that rank collapse is not only an issue during inference but also hinders training due to vanishing gradient problems at initialization. In Wu et al. [2024b] the authors show that rank collapse also happens in graph attention networks (in this setting the phenomenon is also known as oversmoothing), also causing the model to lose expressive power. Finally, the work in Wu et al. [2024a] studies the effect of LayerNorm on rank collapse in transformers, by proving that this component can help preventing or mitigating rank collapse.
@@ 3.1 Sequence Model Deep-Learning Architecture
The main mechanism differs between attention in transformers, and a recurrent block in SSMs. However, as shown in Ali et al. [2024] and Dao and Gu [2024], both models can be seen as part of a general framework
@@
[…]
@@ 4.1 Lambda-Skip Connection: sufficient to prevent rank collapse
$b:=\frac{1}{a^{K}}\frac{2\lambda NdSC_{M}}{\lambda^{2}-a(SC_{M}+|\lambda|)^{2}},$
with $C_{M}:=\sup_{k}||M^{(k)}||_{F}$, $S:=\sup_{k}||C_{V}^{(k)}||_{F}$.We also define the collapse rate as:
@@
The key part of the proof is to show the following relationship between the rank collapse measure at two consecutive layers: $\mu(Y^{(k+1)})^{2}\geq\frac{1}{(SC_{M}+|\lambda|)^{2}}\left(\lambda^{2}\mu(Y^{(k)})^{2}-2\lambda NdC_{M}\right)$, which is done by algebraic manipulation and upper bounds on $||Y^{(k+1)}||_{F}$. The claim follows by imposing the right-hand side above greater of equal to $a\mu(Y^{(k)})^{2}$, rearranging the terms and imposing the condition on $\lambda$ to guarantee the feasibility of this. The full proof can be found in Appendix A.4. ∎
We note that this result is very general, since it holds for any model that can be expressed in the form of equation 6, independent of its specific implementation or parametrization. Moreover, the presence of LayerNorm implies that the bound provided in equation 7 does not depend on the input even when the matrix $M$ is input-dependent. This is because the input is normalized before being fed through $M$. Specifically, $C_{M}$ is only dependent on the weights and on the sequence length  [^2] In transformers $C_{M}=\sqrt{N}$, in structured LTI SSMs $C_{M}\geq\sup_{k}||A^{(k)}||\ ||B^{(k)}||\ ||C^{(k)}||_{F}$ and in selective SSMs $C_{M}\geq\sup_{k}||W_{B}^{(k)}||\ ||W_{C}^{(k)}||_{F}$ (since in Selective SSMs it is common practice to choose $A_{t}^{(k)}$ with eigenvalues smaller than one).
Theorem 4.1 also provides an important intuition for the choice of $\lambda$. In particular, we note that to guarantee prevention of rank collapse, the ideal choice is $a=1$. However, this choice will be mediated by the values of $C_{M}$ and $S$ in the different architectures. In particular, it is easy to see that the only way to guarantee a solution to 7 is by having $01-a>$ 0, for which $\lambda$ needs to satisfy $|\lambda|>\frac{(a+\sqrt{a})SC_{M}}{1-a}$
@@
Finally, we conclude this section by briefly discussing how Theorem 4.1 can be generalized to the case where we let $\lambda$ vary across different layers. The following small modification will be enough to make the result still hold in this setting: in order to still satisfy the condition $\mu(Y^{(k)})\geq a\mu(Y^{(k-1)})$, we can simply adapt equation 7 to be $\lambda_{k}^{2}-a(S_{k}C_{M_{k}}+|\lambda_{k}|)^{2}>0$, where $S_{k}=||C^{(k)}_{V}||_{F}$ and $C_{M_{k}}=||M^{(k)}||_{F}$. Additionally, the flexibility to choose $\lambda_{k}$ allows for different values of $a$ to be used across layers. For instance, instead of setting a single $a$ value for all layers, one can vary $\lambda_{k}$ to meet different layer-specific conditions. In one layer, $\lambda_{k}$ might satisfy the condition for a smaller value of $a$, while in a subsequent layer, $\lambda_{k}$ satisfies the condition for a larger value of $a$.
@@ 4.2.1 Rank collapse occurs without skip connections
We refer to Appendix A.5 for a rigorous presentation of Theorem 4.2. The main takeaway from this result is that, under appropriate conditions on the architecture and the input sequence, removing skip connections and only using attention layers and LayerNorm can cause an exponentially fast convergence to a rank-$1$ matrix. This highlights the importance of skip connections to mitigate rank collapse. It is also interesting to compare the result in Theorem 4.2 to the setting of self-attention only transformers (without both LayerNorm and skip connections), reported in Dong et al. [2023]. Under certain assumptions on the attention matrix and the input sequence, rank collapse occurs at a doubly exponential rate. We report the full result in Theorem A.3 and Corollary A.3.1 (where we adapted the Theorem to the rank collapse measure provided in the problem statement).
@@ 4.2.2 Rank Collapse can still occur with skip connections under specific choices of $\lambda$: examples
In the following, we show an analytical example of how, for Selective SSMs, ank collapse can still occur in the presence of lambda-skip connections for specific choice of $\lambda$. A similar example for Structured LTI SSMs can be found in Appendix A.10 .
@@
Consider the system, which we will denote as [Sys-2], $A_{t}=A=\alpha I$ (for simplicity, we consider $\alpha$ not to be input dependent here), $B=YW_{B}$ and $C=YW_{C}$ input dependent (where $y_{t}$ is the $t$-th row of $Y$) with LayerNorm applied to the output of each layer. Once again, we let $N=2$ and $d=2$. We have $M=W_{A}\odot\left(YW_{C}W_{B}^{\top}Y^{\top}\right)$ where $W_{A}=\begin{pmatrix}1&0\\ \alpha&1\\ \end{pmatrix}$. By choosing $\alpha=1$ and $W_{B}=W_{C}=I$, we get that $M=\begin{pmatrix}1&0\\ y_{1}^{\top}y_{2}&1\end{pmatrix}$ since the input to the layer is normalized by LayerNorm.
@@
4.2.3 Tightness of Lower Bound in Theorem 4.1
We conclude this section with a discussion of the tightness of the lower bound provided in Theorem 4.1. In particular, we explore whether there exist a system such that $\mu(Y^{(k)})^{2}=O\left(a^{k}\mu(Y^{(0)})^{2}\right)$ for a choice of $\lambda$ that satisfies Equation 7. We show in Proposition 4.3.2 that such a system exists. Hence, without additional assumptions on the architecture or on the input sequence, it is not possible to provide a better guarantee for the lower bound of the rank collapse metric.
Proposition 4.3.2.
There exist an architecture of the form of Equation 6 such that, when choosing $\lambda$ to satisfy Equation 7, i.e. $|\lambda|=\Omega(\frac{a}{1-a})$, it holds that $\mu(Y^{(k)})^{2}=O\left(a^{k}\mu(Y^{(0)})^{2}\right)$.
Proof.
We again consider [Sys-2] from Proposition 4.3.1 with $\alpha_{0}=\beta_{0}=\frac{1}{2}$. From the proof of Proposition 4.3.1, we calculate the values of $Y^{(k)}$ and $\mu(Y^{(k)})$ for all layers $k$. Then, we calculate the value of $a$ by simply considering the ratio of the rank collapse measure at two consecutive layers. Then, the result simply follows by upper bounding the value of $a$ and comparing it with the desired bound. The full proof can be found in Appendix A.12. ∎